# Struct2D: A Perception-Guided Framework for Spatial Reasoning in Large Multimodal Models

### v1

@@ 3.2 Zero-shot Analysis of Struct2D Prompting
We construct a subset of 422 QA pairs for evaluation, selected due to API call budgets. As shown in Figure 3, the distribution of question types is consistent with the full benchmark. For our analysis, we generate object marks using both ground-truth 3D annotations and noisy detections (following [30, 62]), ensuring comprehensive object coverage while eliminating perception errors. This also enables a fair comparison with prior work, particularly [62].
Table 1: Zero-shot evaluation of GPT-o3 on the VSI-Bench subset. The first row simply uses 16 frames from the input video, proposed in VSI-Bench [81]. For our prompting, we only input a BEV image with object marks on it along with object-centric meta information.
@@
VSI-Bench [81] 16 105.07 48.6 44.3 34.1 50.9 51.0 49.4 61.9
GPT4Scene [62] 9 78.67 50.3 51.5 35.3 58.0 50.5 47.9 58.8
@@
Zero-shot Prompting Results. Table 1 shows that GPT-o3 exhibits strong spatial reasoning capabilities when prompted with structured 2D inputs. Specifically, providing both object-centric metadata and filtered object marks significantly boosts performance, achieving 96.5 on relative distance, 94.4 on relative direction, and 80.1 on route planning. This highlights that explicit 3D representations are not strictly necessary—LMMs can reason effectively with carefully structured 2D projections. The ablation further reveals that rotation alignment and a structured guide prompt each contribute to improved accuracy on relative direction tasks, with the combination of both yielding the best performance (94.4). These results underscore the importance of aligning spatial context and guiding the model through geometric reasoning steps. Notably, our method requires only a single BEV image and lightweight metadata, making it a low-cost and robust alternative to multi-frame prompting strategies [81, 62].
@@ 4 Large-Scale Instruction Tuning with Struct2D-Set
In this section, we first describe the design and construction of Struct2D-Set, highlighting its coverage, annotation pipeline, and task diversity. We then present the supervised fine-tuning (SFT) setup using open-source LMMs, detailing the model configurations and training procedures. Finally, we evaluate the effectiveness of our instruction-tuned model across multiple spatial reasoning benchmarks, assessing its generalization and reasoning capabilities.
@@ 4.1 Struct2D-Set
Overview. Struct2D-Set consists of 200K QA pairs generated from over 6K richly annotated indoor scenes, sourced from large-scale 3D reconstruction datasets—ARKitScenes[4], ScanNet [19], and ScanNet++[87]. These datasets capture diverse real-world environments, including homes, offices, and industrial settings. The QA pairs cover eight categories of spatial reasoning tasks. Figure 4 shows the distribution of question types across the dataset.
@@
The first type, inspired by VSI-Bench [81], involves tasks that require understanding global spatial relationships in 3D, such as spatial relation identification, egocentric navigation, and comparative reasoning. These questions cannot be answered from a single keyframe alone. We begin by extracting ground-truth object annotations from the training sets of the 3D datasets, including object boxes, depth maps, and camera poses. Using structured templates, we generate initial QA pairs based on this meta information, and then enrich them using ChatGPT to produce step-by-step reasoning traces and more natural language formulations. Each QA pair includes a short answer derived from geometry templates and a long-form answer elaborating on the reasoning process.
The second type of QA pairs is adapted from existing 3D scene understanding benchmarks, including ScanQA [3], SQA3D [52], Scan2Cap [14], ScanRefer [10], and Multi3DRefer [94]. These examples cover tasks such as object attribute identification, counting, and binary verification. We augment the original training set questions and descriptions using ChatGPT to improve clarity and reasoning depth. These tasks typically benefit from selecting keyframes where relevant objects are clearly visible, allowing the model to ground spatial reasoning in egocentric frames.
@@ 4.2 Experiment Setup
We fine-tune the open-source LMM Qwen2.5VL [71] using our proposed dataset, Struct2D-Set. For evaluation, we primarily focus on VSI-Bench [81], which includes complex spatial reasoning tasks such as relative direction and route planning. Additionally, we assess model performance on three standard 3D scene understanding tasks built on ScanNet [19]: 3D question answering (ScanQA [3], SQA3D [52]), 3D dense captioning (Scan2Cap [14]), and 3D visual grounding (ScanRefer [10], Multi3DRef [94]). For VSI-Bench, we input only BEV images with filtered object marks and metadata, as the tasks focus purely on spatial relationships. For the other benchmarks, which often involve object attributes or visual details, we additionally provide selected egocentric keyframes to support reasoning.
@@ 4.3 Implementation Details
We adopt Qwen2.5VL [71] as our base LMM for instruction tuning. During training, the model receives BEV images with filtered object marks and object-centric metadata. For tasks that require appearance or attribute information (e.g., object color or count), we additionally provide egocentric keyframes. All visual inputs are resized to $480\times 480$, and object marks are adaptively scaled based on their original resolution.
For questions involving complex spatial reasoning, such as relative direction or route planning, we insert special tokens <think> and </think> to guide the model to generate a step-by-step reasoning process, followed by the final answer enclosed within <answer> and </answer>. For simpler questions involving object appearance or quantitative estimation, the model is trained to directly produce short answers without reasoning traces. We train the model for one epoch using a base learning rate of 2e-6 with cosine annealing. Training takes approximately 8 hours on 4×H200 GPUs. For evaluation, we follow [30, 62] by reconstructing point clouds offline using BundleFusion [18], detecting 3D object boxes with Mask3D and UniDet, and projecting them into BEV images and 2D object marks.
Table 3: Performance comparison of various models on VSI-Bench [81]. The model fine-tuned with Struct2D-Set surpasses both the Struct2D prompting and the video-based tuning baseline.
@@
InternVL2-2B [15] 30.3 21.8 24.9 35.0 22.0 33.8 44.2 30.5
InternVL2-8B [15] 33.9 23.1 28.7 39.8 48.2 36.7 30.7 29.9
LongVILA-8B [79] 21.1 29.1 9.1 0.0 16.7 29.6 30.7 32.5
VILA-1.5-8B [45] 29.5 17.4 21.8 18.8 50.3 32.1 34.8 31.0
LongVA-7B [90] 31.1 38.0 16.6 22.2 38.9 33.1 43.3 25.4
LLaVA-NeXT-Video-7B [95] 36.3 48.5 14.0 24.2 47.8 43.5 42.4 34.0
LLaVA-OneVision-0.5B [39] 31.2 46.1 28.4 28.3 15.4 28.9 36.9 34.5
LLaVA-OneVision-7B [39] 33.5 47.7 20.2 12.3 47.4 42.5 35.2 29.4
R1-Zero-VSI [44] (Qwen2-VL-7B) 32.1 39.4 25.0 43.2 25.8 32.6 30.9 27.8
R1-Zero-VSI [44] (Qwen2-VL-7B) + SFT 38.8 44.7 27.6 50.4 46.1 34.0 35.7 33.0
@@
Table 4: 3D Question Answering Evaluation on ScanQA [3] and SQA3D [52] datasets.
@@
ScanQA [3] 30.2 10.1 13.1 33.3 64.9
SQ3D [52] 46.6
3D-VLP [33] 30.5 11.2 13.5 34.5
3D-Vista [101] 13.9 35.7 48.5
@@
Chat-3D [72] 29.1 6.4 11.9 28.5 53.2
Chat-3D v2 [30] 38.4 7.3 16.1 40.1 77.1
3D-LLM [27] 39.3 12.0 14.5 37.3 69.4
LL3DA [13] 13.5 15.9 37.3 76.8
PQ3D [102] 47.1
LEO [31] 11.5 16.2 39.3 80.0 50.0 50.0
Chat-Scene [30] 43.2 14.3 18.0 41.6 87.7 54.6 57.5
@@
InternVL-2-8B [15] 23.9 3.3 14.5 34.3 62.5 33.0 45.3
MiniCPM-V-2.6 [86] 25.1 8.4 11.8 31.5 60.1 42.6 46.6
@@
4.4 Main results
We present quantitative results on VSI-Bench[81] in Table 3 and on ScanQA[3] and SQA3 [52] in Table 4. Additional benchmark results are provided in the Appendix due to space limitations.
As shown in Table 3, our model fine-tuned with the Struct2D-Set dataset achieves the highest average score (41.9) among all open-source models evaluated on VSI-Bench. Notably, it surpasses both the Struct2D prompting variant (29.4) and the standard baseline trained with uniformly sampled 16 video frames (33.9), confirming the effectiveness of our full instruction tuning approach. The performance gains are especially prominent on spatial reasoning tasks like relative direction (44.9) and route planning (33.5), where the model must integrate both geometric understanding and egocentric context. Compared with R1-Zero-VSI [44] (38.8), a recent method that trains Qwen2-VL-7B using video-based supervision, our tuned model not only achieves stronger average performance but also uses fewer visual frames and does not rely on dense temporal input. These results highlight the scalability and efficiency of Struct2D-Set for training capable spatial reasoners without explicit 3D features at inference.
Table 4 shows results on two traditional 3D question answering benchmarks, ScanQA and SQA3D. Our model outperforms most existing methods, including several that rely on explicit 3D point cloud inputs. Compared with GPT4Scene [62], our model performs on par across most metrics. However, these benchmarks primarily require identifying relevant keyframes and generating free-form textual answers. As a result, models can often rely on memorizing object-level attributes, and the rule-based evaluation metrics (e.g., BLEU, CIDEr) may not fully reflect the correctness or reasoning quality of the generated answers. Please refer to Appendix for qualitative results.
@@ 6 Conclusion
We present Struct2D, a perception-guided framework that enables LMMs to perform 3D spatial reasoning using structured 2D inputs. Through zero-shot analysis and instruction tuning, we show that BEV images, object-centric metadata, and keyframes are sufficient to unlock strong spatial reasoning capabilities—without requiring explicit 3D inputs. Our curated dataset, Struct2D-Set, supports scalable instruction tuning with fine-grained QA pairs grounded in real 3D scenes. Fine-tuning with Struct2D-Set yields significant gains across spatial reasoning benchmarks, outperforming prior open-source methods even under noisy perception. These findings demonstrate that structured 2D projections are a practical and effective alternative to direct 3D representations, offering a scalable path toward robust multimodal spatial understanding in LMMs.
@@
To select keyframes, we sample $N$ RGB-D frames and iteratively check for visibility of objects not yet covered in the BEV. For each candidate frame $I_{i}$, we project the remaining unseen objects onto both the frame and its depth map. If a valid projection exists (i.e., the projected location lies within the image and has valid depth), the object mark is rendered and the frame is added to the keyframe set $\mathcal{I}_{\text{keys}}$. This process continues until all relevant objects are covered. The final prompt consists of ❶ a BEV image with filtered marks, ❷ optional keyframes containing visible objects, and ❸ object metadata text, all of which are passed to a large multimodal model for reasoning.
This framework allows the LMM to perform 3D spatial reasoning from 2D visual and textual inputs, without requiring direct access to raw 3D data at inference time. It enables scalable, flexible spatial understanding grounded in realistic perception outputs.
@@
7:Sample $N$ frames: $\mathbf{I},\mathbf{D_{I}}\leftarrow\text{Sample}(\mathbf{V},\mathbf{D})$
@@
Figure 7: Overview of the Struct2D Prompting Framework. Given an egocentric video and a spatial question, we first reconstruct a 3D point cloud and remove the ceiling for a clear top-down view. Objects are detected in 3D space, and a bird’s-eye-view (BEV) image is rendered with object marks projected onto the floor plane. These object marks are filtered based on the content of the question. We also extract egocentric keyframes by projecting 3D bounding box centers onto the video, when appearance cues are needed. Object-centric metadata—including object IDs and 3D coordinates—is encoded as text. The structured 2D visual and textual inputs are then fed into a large multimodal model for spatial reasoning.
Qualitative comparison of our Struct2D prompting. To better understand the impact of prompt design on spatial reasoning, we conduct qualitative analyses highlighting two key aspects of our framework: reasoning guidance, object orientation, and structured metadata. As shown in Fig. 8, when the model is prompted only with a BEV image and object marks, it struggles to accurately resolve relative spatial relationships. Adding a structured guide prompt enables the model to decompose the task into interpretable geometric steps, though it may still fail without an aligned reference frame. Once the BEV is rotation-aligned with the agent’s viewpoint, the reasoning becomes more intuitive, leading to the correct answer. Similarly, in Fig. 9, we illustrate the benefit of object-centric metadata. Without access to precise coordinates, the model must estimate distances visually, which can lead to errors. When provided with 3D object positions, the model can directly compute spatial relations such as Euclidean distances, significantly improving its accuracy on localization tasks. These examples highlight how prompt structure—through guided reasoning and geometric priors—plays a crucial role in unlocking spatial understanding in LMMs.
Figure 8: Effects of guide prompting and rotation correction in relative spatial reasoning of the zero-shot prompting. We show different variations of our Struct2D. The original model struggles to interpret spatial relationships using the raw BEV image and object marks alone, failing to answer correctly. With the addition of a structured guide prompt, the model is able to decompose the reasoning process into geometric steps, though it still arrives at an incorrect answer. When object orientation is further corrected via rotation alignment, the spatial reference frame becomes more intuitive, and the model’s reasoning becomes clearer and more accurate. This demonstrates the importance of both prompting design and canonical alignment in unlocking LMMs’ spatial understanding abilities. (Red texts are wrong answers; Green texts are correct ones.)

### v3

@@ 3.2 Zero-shot Analysis of Struct2D Prompting
We construct a subset of 422 QA pairs for evaluation, selected due to API call budgets. As shown in Figure 3, the distribution of question types is consistent with the full benchmark. For our analysis, we generate object marks using both ground-truth 3D annotations and noisy detections (following [huang2023chat, qi2025gpt4scene]), ensuring comprehensive object coverage while eliminating perception errors. This also enables a fair comparison with prior work, particularly [qi2025gpt4scene].
Table 1: Zero-shot evaluation of GPT-o3 on the VSI-Bench subset. The first row simply uses 16 frames from the input video, proposed in VSI-Bench [yang2024thinking]. For our prompting, we only input a BEV image with object marks on it along with object-centric meta information.
@@
VSI-Bench [yang2024thinking] 16 105.07 48.6 44.3 34.1 50.9 51.0 49.4 61.9
GPT4Scene [qi2025gpt4scene] 9 78.67 50.3 51.5 35.3 58.0 50.5 47.9 58.8
@@
Zero-shot Prompting Results. Table 1 shows that GPT-o3 exhibits strong spatial reasoning capabilities when prompted with structured 2D inputs. Specifically, providing both object-centric metadata and filtered object marks significantly boosts performance, achieving 96.5 on relative distance, 94.4 on relative direction, and 80.1 on route planning. This highlights that explicit 3D representations are not strictly necessary—MLLMs can reason effectively with carefully structured 2D projections. The ablation further reveals that rotation alignment and a structured guide prompt each contribute to improved accuracy on relative direction tasks, with the combination of both yielding the best performance (94.4). These results underscore the importance of aligning spatial context and guiding the model through geometric reasoning steps. Notably, our method requires only a single BEV image and lightweight metadata, making it a low-cost and robust alternative to multi-frame prompting strategies [yang2024thinking, qi2025gpt4scene].
@@ 4 Large-Scale Instruction Tuning with Struct2D-Set
In this section, we first describe the design and construction of Struct2D-Set, highlighting its coverage, annotation pipeline, and task diversity. We then present the supervised fine-tuning (SFT) setup using open-source MLLMs, detailing the model configurations and training procedures. Finally, we evaluate the effectiveness of our instruction-tuned model across multiple spatial reasoning benchmarks, assessing its generalization and reasoning capabilities.
@@ 4.1 Struct2D-Set
Overview. Struct2D-Set consists of 200K QA pairs generated from over 6K richly annotated indoor scenes, sourced from large-scale 3D reconstruction datasets—ARKitScenes[baruch2021arkitscenes], ScanNet [dai2017scannet], and ScanNet++[yeshwanth2023scannet++]. These datasets capture diverse real-world environments, including homes, offices, and industrial settings. The QA pairs cover eight categories of spatial reasoning tasks. Figure 4 shows the distribution of question types across the dataset.
@@
The first type, inspired by VSI-Bench [yang2024thinking], involves tasks that require understanding global spatial relationships in 3D, such as spatial relation identification, egocentric navigation, and comparative reasoning. These questions cannot be answered from a single keyframe alone. We begin by extracting ground-truth object annotations from the training sets of the 3D datasets, including object boxes, depth maps, and camera poses. Using structured templates, we generate initial QA pairs based on this meta information, and then enrich them using ChatGPT to produce step-by-step reasoning traces and more natural language formulations. Each QA pair includes a short answer derived from geometry templates and a long-form answer elaborating on the reasoning process.
The second type of QA pairs is adapted from existing 3D scene understanding benchmarks, including ScanQA [azuma2022scanqa], SQA3D [ma2022sqa3d], Scan2Cap [chen2021scan2cap], ScanRefer [chen2020scanrefer], and Multi3DRefer [Multi3DRefer]. These examples cover tasks such as object attribute identification, counting, and binary verification. We augment the original training set questions and descriptions using ChatGPT to improve clarity and reasoning depth. These tasks typically benefit from selecting keyframes where relevant objects are clearly visible, allowing the model to ground spatial reasoning in egocentric frames.
@@ 4.2 Experiment Setup
We fine-tune the open-source MLLM Qwen2.5VL [wang2024qwen2] using our proposed dataset, Struct2D-Set. For evaluation, we primarily focus on VSI-Bench [yang2024thinking], which includes complex spatial reasoning tasks such as relative direction and route planning. Additionally, we assess model performance on three standard 3D scene understanding tasks built on ScanNet [dai2017scannet]: 3D question answering (ScanQA [azuma2022scanqa], SQA3D [ma2022sqa3d]), 3D dense captioning (Scan2Cap [chen2021scan2cap]), and 3D visual grounding (ScanRefer [chen2020scanrefer], Multi3DRef [Multi3DRefer]). For VSI-Bench, we input only BEV images with filtered object marks and metadata, as the tasks focus purely on spatial relationships. For the other benchmarks, which often involve object attributes or visual details, we additionally provide selected egocentric keyframes to support reasoning.
@@ 4.3 Implementation Details
We adopt Qwen2.5VL [wang2024qwen2] as our base MLLM for instruction tuning. During training, the model receives BEV images with filtered object marks and object-centric metadata. For tasks that require appearance or attribute information (e.g., object color or count), we additionally provide egocentric keyframes. All visual inputs are resized to $480\times 480$, and object marks are adaptively scaled based on their original resolution.
For questions involving complex spatial reasoning, such as relative direction or route planning, we insert special tokens <think> and </think> to guide the model to generate a step-by-step reasoning process, followed by the final answer enclosed within <answer> and </answer>. For simpler questions involving object appearance or quantitative estimation, the model is trained to directly produce short answers without reasoning traces. We train the model for one epoch using a base learning rate of 2e-6 with cosine annealing. Training with the whole Struct2D-Set takes approximately 8 hours on 8×H200 GPUs. For evaluation, we follow [huang2023chat, qi2025gpt4scene] by reconstructing point clouds offline using BundleFusion [dai2017bundlefusion], detecting 3D object boxes with Mask3D and UniDet, and projecting them into BEV images and 2D object marks.
Table 3: Performance comparison of various models on VSI-Bench [yang2024thinking]. The model fine-tuned with Struct2D-Set surpasses both the Struct2D prompting and the video-based tuning baseline.
@@
InternVL2-2B [chen2024internvl] 30.3 21.8 24.9 35.0 22.0 33.8 44.2 30.5
InternVL2-8B [chen2024internvl] 33.9 23.1 28.7 39.8 48.2 36.7 30.7 29.9
LongVILA-8B [xue2024longvila] 21.1 29.1 9.1 0.0 16.7 29.6 30.7 32.5
VILA-1.5-8B [lin2023vila] 29.5 17.4 21.8 18.8 50.3 32.1 34.8 31.0
LongVA-7B [zhang2024longva] 31.1 38.0 16.6 22.2 38.9 33.1 43.3 25.4
LLaVA-NeXT-Video-7B [zhang2024llavanext-video] 36.3 48.5 14.0 24.2 47.8 43.5 42.4 34.0
LLaVA-OneVision-0.5B [li2024llava] 31.2 46.1 28.4 28.3 15.4 28.9 36.9 34.5
LLaVA-OneVision-7B [li2024llava] 33.5 47.7 20.2 12.3 47.4 42.5 35.2 29.4
R1-Zero-VSI [liao2025improved] (Qwen2-VL-7B) 32.1 39.4 25.0 43.2 25.8 32.6 30.9 27.8
R1-Zero-VSI [liao2025improved] (Qwen2-VL-7B) + SFT 38.8 44.7 27.6 50.4 46.1 34.0 35.7 33.0
@@
Qwen2.5-VL-7B (SFT) 43.6 47.1 35.1 48.9 57.1 35.1 45.9 35.8
Table 4: 3D Question Answering Evaluation on ScanQA [azuma2022scanqa] and SQA3D [ma2022sqa3d] datasets.
@@
ScanQA [azuma2022scanqa] 30.2 10.1 13.1 33.3 64.9
SQ3D [ma2022sqa3d] 46.6
3D-VLP [jin2023context] 30.5 11.2 13.5 34.5
3D-Vista [zhu20233d] 13.9 35.7 48.5
@@
Chat-3D [wang2023chat] 29.1 6.4 11.9 28.5 53.2
Chat-3D v2 [huang2023chat] 38.4 7.3 16.1 40.1 77.1
3D-LLM [hong20233d] 39.3 12.0 14.5 37.3 69.4
LL3DA [chen2024ll3da] 13.5 15.9 37.3 76.8
PQ3D [zhu2024unifying] 47.1
LEO [huang2023embodied] 11.5 16.2 39.3 80.0 50.0 50.0
Chat-Scene [huang2023chat] 43.2 14.3 18.0 41.6 87.7 54.6 57.5
@@
InternVL-2-8B [chen2024internvl] 23.9 3.3 14.5 34.3 62.5 33.0 45.3
MiniCPM-V-2.6 [yao2024minicpm] 25.1 8.4 11.8 31.5 60.1 42.6 46.6
@@
4.4 Main results
We present quantitative results on VSI-Bench[yang2024thinking] in Table 3 and on ScanQA[azuma2022scanqa] and SQA3 [ma2022sqa3d] in Table 4. Additional benchmark results are provided in the Appendix due to space limitations.
As shown in Table 3, our model fine-tuned with the Struct2D-Set dataset achieves the highest average score (43.6) among all open-source models evaluated on VSI-Bench. Notably, it surpasses both the Struct2D prompting variant (29.4) and the standard baseline trained with uniformly sampled 16 video frames (33.9), confirming the effectiveness of our full instruction tuning approach. The performance gains are especially prominent on spatial reasoning tasks like relative direction (45.9) and route planning (35.8), where the model must integrate both geometric understanding and egocentric context. Compared with R1-Zero-VSI [liao2025improved] (38.8), a recent method that trains Qwen2-VL-7B using video-based supervision, our tuned model not only achieves stronger average performance but also uses fewer visual frames and does not rely on dense temporal input. These results highlight the scalability and efficiency of Struct2D-Set for training capable spatial reasoners without explicit 3D features at inference.
Table 4 shows results on two traditional 3D question answering benchmarks, ScanQA and SQA3D. Our model outperforms most existing methods, including several that rely on explicit 3D point cloud inputs. Compared with GPT4Scene [qi2025gpt4scene], our model performs on par across most metrics. However, these benchmarks primarily require identifying relevant keyframes and generating free-form textual answers. As a result, models can often rely on memorizing object-level attributes, and the rule-based evaluation metrics (e.g., BLEU, CIDEr) may not fully reflect the correctness or reasoning quality of the generated answers. Please refer to Appendix for qualitative results.
@@ 6 Conclusion
We present Struct2D, a perception-guided framework that enables MLLMs to perform 3D spatial reasoning using structured 2D inputs. Through zero-shot analysis and instruction tuning, we show that BEV images, object-centric metadata, and keyframes are sufficient to unlock strong spatial reasoning capabilities—without requiring explicit 3D inputs. Our curated dataset, Struct2D-Set, supports scalable instruction tuning with fine-grained QA pairs grounded in real 3D scenes. Fine-tuning with Struct2D-Set yields significant gains across spatial reasoning benchmarks, outperforming prior open-source methods even under noisy perception. These findings demonstrate that structured 2D projections are a practical and effective alternative to direct 3D representations, offering a scalable path toward robust multimodal spatial understanding in MLLMs.
Acknowledgment
We thank Zhangyang Qi for the thoughtful discussions.
@@ Acknowledgment
To select keyframes, we sample $N$ RGB-D frames and iteratively check for visibility of objects not yet covered in the BEV. For each candidate frame $I_{i}$, we project the remaining unseen objects onto both the frame and its depth map. If a valid projection exists (i.e., the projected location lies within the image and has valid depth), the object mark is rendered and the frame is added to the keyframe set $\mathcal{I}_{\text{keys}}$. This process continues until all relevant objects are covered. The final prompt consists of ❶ a BEV image with filtered marks, ❷ optional keyframes containing visible objects, and ❸ object metadata text, all of which are passed to a multimodal large language model for reasoning.
This framework allows the MLLM to perform 3D spatial reasoning from 2D visual and textual inputs, without requiring direct access to raw 3D data at inference time. It enables scalable, flexible spatial understanding grounded in realistic perception outputs.
@@
7:Sample NN frames: $\mathbf{I},\mathbf{D_{I}}\leftarrow\text{Sample}(\mathbf{V},\mathbf{D})$
@@
Figure 7: Overview of the Struct2D Prompting Framework. Given an egocentric video and a spatial question, we first reconstruct a 3D point cloud and remove the ceiling for a clear top-down view. Objects are detected in 3D space, and a bird’s-eye-view (BEV) image is rendered with object marks projected onto the floor plane. These object marks are filtered based on the content of the question. We also extract egocentric keyframes by projecting 3D bounding box centers onto the video, when appearance cues are needed. Object-centric metadata—including object IDs and 3D coordinates—is encoded as text. The structured 2D visual and textual inputs are then fed into a multimodal large language model for spatial reasoning.
Qualitative comparison of our Struct2D prompting. To better understand the impact of prompt design on spatial reasoning, we conduct qualitative analyses highlighting two key aspects of our framework: reasoning guidance, object orientation, and structured metadata. As shown in Figure 8, when the model is prompted only with a BEV image and object marks, it struggles to accurately resolve relative spatial relationships. Adding a structured guide prompt enables the model to decompose the task into interpretable geometric steps, though it may still fail without an aligned reference frame. Once the BEV is rotation-aligned with the agent’s viewpoint, the reasoning becomes more intuitive, leading to the correct answer. Similarly, in Figure 9, we illustrate the benefit of object-centric metadata. Without access to precise coordinates, the model must estimate distances visually, which can lead to errors. When provided with 3D object positions, the model can directly compute spatial relations such as Euclidean distances, significantly improving its accuracy on localization tasks. These examples highlight how prompt structure—through guided reasoning and geometric priors—plays a crucial role in unlocking spatial understanding in MLLMs.
Figure 8: Effects of guide prompting and rotation correction in relative spatial reasoning of the zero-shot prompting. We show different variations of our Struct2D. The original model struggles to interpret spatial relationships using the raw BEV image and object marks alone, failing to answer correctly. With the addition of a structured guide prompt, the model is able to decompose the reasoning process into geometric steps, though it still arrives at an incorrect answer. When object orientation is further corrected via rotation alignment, the spatial reference frame becomes more intuitive, and the model’s reasoning becomes clearer and more accurate. This demonstrates the importance of both prompting design and canonical alignment in unlocking MLLMs’ spatial understanding abilities. (Red texts are wrong answers; Green texts are correct ones.)